Let A and B be independent events with pr(A)= .3 and pr(B)b=.4 What is the probability that exactly one of the events A or B occurs?
Answer by ikleyn(53751)

.

The probability that at least one of the events A or B occurs is equal to the sum of probabilities P(A) + P(B) minus 
the probability that BOTH events (A & B) occurs at the same time:


    P( at least one of the events A or B occurs ) = P(A) + P(B) - P(A & B).


Since the events A and B are independent,  P(A & B) = P(A)*P(B) = 0.3*0.4 = 0.12,  so  the probability 
that at least one of the two events will happen is


    P( at least one of the two events will happens ) = P(A) + P(B) - P(A)*P(B) = 0.3 + 0.4 - 0.3*0.4 = 0.58.   


From this, we should subtract the probability that BOTH events (A & B) occur.


So, the final answer to the problem question is


       P( exactly one of the events A or B occurs ) = P(A) + P(B) - 2P(A & B) = 0.58 - 0.12 = 0.46.       ANSWER


Another way to derive the same formula and to get the same answer is THIS :

    the value  P(A) - P(A & B)  is the probability that only A will occur with no B;

    the value  P(B) - P(A & B)  is the probability that only B will occur with no A.


    Therefore, the sum  P(A) + P(B) - 2*P(A & B)  is the probability under the question.